Air contains nearly 20%of oxygen by volume. the volume of air needed for complete combustion of acetylene will be....

Grade:12th pass

2 Answers

Arun
5 years ago

2C2H2 + 5O2 → 4CO2 + 2H2O

2 moles of C2H2 react with 5 moles of O2

22.44 L of C2H2 reacts with (5/2) x 22.2 L moles of O2

Hence 100 cc of C2H2 will react with (5/2)x 100 cc = 250 cc of O2

Since air contains 20% of oxygen by volume, the amount of air needed to react with 100 cc of C2H2 = 250 x (100/20) = 1250 cc of air

DONA KURIAN
3 years ago
2C2H2 + 5O2 → 4 CO2 + 2H2O
2 mole ->5 mole
2mL         5mL
100mL      z x 20/100                                 cross multiply
then,
        z x 40/100 =500
     
         therefore ,z = 1250